OpenStudy (anonymous):

Solve for x. 6 – 3(2x – 4) = 2(1 + x) A. –2 B. 0 C. –1 D. 2

OpenStudy (anonymous):

Start by distributing

OpenStudy (anonymous):

Thats my problem i dont know if to distribute -3 or just +3 to )2x-4)

OpenStudy (anonymous):

@ASNinja

OpenStudy (anonymous):

it'd be the -3

OpenStudy (anonymous):

ok what u do is 6x-12=2+x

OpenStudy (anonymous):

so now what we do is move the x in one place and the numbers in one so it would be -14=-5x

OpenStudy (anonymous):

Oh okay thanks

OpenStudy (texaschic101):

hzer...don't forget the 6 at the beginning of the problem

OpenStudy (anonymous):

I get 2.8 as an answer, but it isnt there. @hzer

OpenStudy (texaschic101):

6 - 3(2x - 4) = 2(1 + x) --- distribute through the parenthesis 6 - 6x + 12 = 2 + 2x -6x + 18 = 2 + 2x --- subtract 2x from both sides -6x - 2x + 18 = 2 -- subtract 18 from both sides -6x - 2x = 2 - 18 -- combine like terms -8x = - 16 -- divide both sides by -8 x = -16/-8 x = 2

OpenStudy (texaschic101):

any questions ?